# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ Abstract
Differential privacy is a well-established framework for safeguarding sensitive information in data. While extensively applied across various domains, its application to network data — particularly at the node level — remains underexplored. Existing methods for node-level privacy either focus exclusively on query-based approaches, which restrict output to pre-specified network statistics, or fail to preserve key structural properties of the network. In this work, we propose GRAND (Graph Release with Assured Node Differential privacy), which is, to the best of our knowledge, the first network release mechanism that releases entire networks while ensuring node-level differential privacy and preserving structural properties. Under a broad class of latent space models, we show that the released network asymptotically follows the same distribution as the original network. The effectiveness of the approach is evaluated through extensive experiments on both synthetic and real-world datasets.
@@ 1 Introduction
Rapid technological advances and explosive social connectivity have made complex networks more common than ever. From the Internet to transportation topologies for traffic planning and international trade agreements, networks are widely used to depict intensive interactions within complex systems. However, nestled within these marvels of modernity are layers of privacy concerns that cannot be overlooked. For example, in studies of organ donation networks, many of the donors and recipients are not willing to disclose their identities (Hizo-Abes et al., 2010; Marcus et al., 2023). In studying financial system crisis, firms are generally reluctant to share records of their transactions with other firms. Similarly, in studying the sexual relationship between HIV patients, the identity of subjects has to be protected (Little et al., 2014; Abadie et al., 2021).
These concerns manifest at various levels – ranging from personal specter of private information leakage, at the macroscopic level, affecting entire donation or financial relations with the potential to precipitate systemic crises. At the forefront are safeguarding or preventive actions that intend to avoid potential risks preemptively, which focus on the fortification of complex network data against privacy breaches. In comparison to conventional (non-network) data-sharing scenarios, enhancing privacy within complex networks presents unique difficulties. These mainly arise from network intricate structures, the prevalent noise, and the vast scale of associated issues, which complicate the use of simple and well-established mathematical models for privacy protection (Karwa et al., 2017; Hehir et al., 2022). Further, algorithms are usually applied to large-scale networks, including social networks on the Internet or computer clusters; these complex networks carry higher requirements regarding algorithms’ computational and communication efficiency (Guo et al., 2023).
@@
Differential privacy (DP; (Dwork, 2006)) has recently become a common practice for privacy protection. It intends to enforce strict privacy protection for all individuals in a dataset, including those with extreme values that are the most vulnerable to privacy leakages and deanonymization. A variety of privacy-preserving mechanisms have been proposed to guarantee differential privacy, such as the Laplace mechanism (Dwork et al., 2006, 2014) and the exponential mechanism (McSherry and Talwar, 2007). These methods have been widely applied in various learning tasks (Lei et al., 2018; Soto et al., 2022; Lin et al., 2023; Xue et al., 2024; Ma et al., 2025). And variations of differential privacy have also been thoroughly studied, including relaxed or approximate differential privacy (Abadi et al., 2016; Cai et al., 2019), local differential privacy (Rohde and Steinberger, 2018), random differential privacy (Hall et al., 2012), Kullback-Leibler privacy (Wang et al., 2016) and Gaussian differential privacy (Dong et al., 2022). Meanwhile, differential privacy has been applied across many fields, both in academia and across industries (Kaissis et al., 2020; Hie et al., 2018; Han et al., 2020; Santos-Lozada et al., 2020; Kenthapadi and Tran, 2018; Nayak, 2020).
@@
Differential privacy in complex network data analysis has also been carefully explored (e.g., Karwa and Slavković, 2016; Rohde and Steinberger, 2018; Chang et al., 2024). However, as network data typically show structures distinct from that of Euclidean data, the preservation of differential privacy for network data entails unique challenges. First, differential privacy preservation may come at the cost of altering a network’s underlying utilities. This includes, but not limited to, crucial network properties, such as degree distributions, centrality and potentially community structures. Many existing methods may inadvertently alter the underlying distribution of edges and degrees. Meanwhile, edge-level modifications, such as edge addition or deletion, may also change node-level properties. In order to preserve a certain network property, therefore, many existing methods focus on statistic release rather than network release – that is, the release of certain summary statistics (e.g., degree distribution or the number of triangles) with privacy guarantee, rather than an entire differentially private network that has all properties preserved.
@@
Second, the adoption of differential privacy may appear at different levels. On the one hand, many existing works consider the concept of edge differential privacy (Nissim et al., 2007), which protects privacy of individual edges. The adoption of edge differential privacy is natural and shows good theoretical properties (Mülle et al., 2015; Karwa et al., 2017; Fan et al., 2020; Hehir et al., 2022; Chang et al., 2024). On the other hand, the investigation and development of node differential privacy (Hay et al., 2009) is relatively rare. By definition, node differential privacy guarantees that a network (or its certain properties) should remain robust with the change of any node. However, the alteration of a node usually entails the alteration of all associated edges (and potentially other information). The preservation of network properties given node differential privacy is therefore considered substantially more challenging (Kasiviswanathan et al., 2013; Hehir et al., 2022). In particular, the release of entire networks that follow node differential privacy has rarely been investigated in previous literature.
@@
In this work, we propose a novel mechanism, GRAND (Graph Release with Assured Node Differential privacy), to release networks that guarantees node differential privacy. Moreover, the proposed method is shown to preserve statistical network properties under the general latent space models, which include many widely used statistical network models as special cases. To the best of our knowledge, this is the first method that can release entire networks (rather than a summary statistic) with node differential privacy that is computationally feasible with utility preservation guarantees.
The rest of this article is organized as follows. Section 2 reviews the background of differential privacy in network data, identifies gaps in the existing literature, and outlines our primary contributions. Section 3 formally defines node differential privacy and also the statistical model class that motivates our design of the GRAND mechanism. Section 4 presents a detailed description of our method. The theoretical properties of GRAND is studied in Section 5. In Section 6, we evaluate GRAND’s ability to preserve network properties through extensive simulation examples. Section 7 further showcases the application of our method in privatizing real-world social networks. We conclude the paper with additional discussions in Section 8.
2 Differential Privacy of Network Data and Our Contributions
Node differential privacy vs. edge differential privacy
@@ 2 Differential Privacy of Network Data and Our Contributions
Broadly speaking, existing works on network differential privacy can be categorized into two branches (Abawajy et al., 2016; Li et al., 2023c). One branch is edge differential privacy (EDP; Nissim et al., 2007). Intuitively, an EDP mechanism ensures that one cannot infer the pairwise relation between any two nodes in the network. Many methods have been proposed and discussed regarding the protection of EDP, which protect the privacy of individual edges. This can be achieved through mechanisms such as randomized responses (Mülle et al., 2015; Hehir et al., 2022), method of moments (Chang et al., 2024), or noise injection (Karwa and Slavković, 2016; Fan et al., 2020). The other branch is node differential privacy (NDP; Hay et al., 2009). An NDP mechanism (to be formally defined later) ensures that one cannot infer an individual node in the network.
@@
Notice that, information associated with a node includes its pairwise relations to other nodes, and hence node differential privacy is more strict than edge differential privacy. As a result, node-level privacy protection is considerably more challenging than edge-level privacy. Applying edge-level DP methods to node-level settings usually leads to degenerate results (Kasiviswanathan et al., 2013; Hehir et al., 2022; Chang et al., 2024).
@@
However, we advocate that preserving node differential privacy can be critical in many scenarios. This is because nodes can represent individual users, customers, and patients, whose identities are sensitive and vulnerable to de-anonymization. Encoding node differential privacy, while preserving the underlying network’s properties is the key motivation of this paper.
Network release vs. statistic release
@@
The adoption of differential privacy in network analysis, either at the node level or at the edge level, involves introducing random noise, which may lead to unintended consequences. First, existing differentially private mechanisms that are designed for Euclidean data may inadvertently change the underlying structure or properties of networks. For example, adding or deleting edges may lead to the alteration of the degree distribution, transitivity, and the number of components. In general, releasing a network that is differentially private while preserving all original properties can be very challenging.
@@
With the aforementioned limitations, it is then evident that directly releasing a network would be substantially more flexible for subsequent use. Ideally, if the released network is almost the same as the original network, known as an informative release (Wasserman and Zhou, 2010), we can use the privatized network in an arbitrary way in downstream tasks. In particular, when the released network has most network properties maintained the same, it essentially subsumes most statistic release. Along this path, many have explored the novel ways of generating synthetic networks under edge differential privacy such as Karwa et al. (2017); Qin et al. (2017); Hehir et al. (2022); Guo et al. (2023); Chang et al. (2024). In contrast, for node differential privacy, the only known methods in the literature (Borgs et al., 2015, 2018; Chen et al., 2024) are NP-hard, thus are only of theoretical interest.
Our contributions
The preceding discussion motivates the need for a mechanism to release entire networks (rather than statistics) with node-level differential privacy, which is expected to be a crucial tool in practice. while such a tool is not yet available in the literature. In this paper, we will introduce a computationally feasible method to release a network under node differential privacy, while preserving key network properties. To the best of our knowledge, we are the first to achieve this goal in the literature.
@@ 3 Node Differential Privacy and Latent Space Models
3.1 Notations
For any positive integer $K$, we denote the set $\{1,\cdots,K\}$ by $[K]$. In particular, we use $[N]$ to index the nodes in the network. Consider a network of $N$ individuals (nodes), where the edges represent interactions or relationships between nodes, such as friendship, collaboration, or transactions (Harris, 2009; Ji and Jin, 2016; Newman, 2018). We assume the network is unweighted and undirected. The network can be uniquely represented by an $N\times N$ adjacency matrix $A$ where $A_{ij}=1$ if individuals $i$ and $j$ are connected, and $A_{ij}=0$ otherwise. Denote the node and edge sets of network $A$ by $\mathcal{V}(A)$ and $\mathcal{E}(A)$, respectively. [^1] The proposed work can be applied to weighted or directed networks as well, as long as a general latent space model defined in Section 3.3 is applicable to the corresponding adjacency matrices. Details will be provided in Section 4.2. For any vector $z\in{\bf R}^{d}$, we use $\|{z}\|$ to denote its Euclidean norm.
@@ 3.2 Node Differential Privacy
Suppose that the network information is sensitive (e.g., a cancer patient network) and many nodes in $A$ may not be willing to release their identity or even their presence in the network. Private information may still be vulnerable to leakage if one simply releases the network to the public, even with de-anonymization (Narayanan and Shmatikov, 2008). In many cases, one may infer a node’s identity according to certain connectivity patterns of a de-anonymized network, such as when one individual is connected with all the rest of the individuals. Alternatively, a node’s identity may be revealed due to adversarial attacks.
We adopt differential privacy (Dwork, 2006) as the standard for privacy protection. One commonly used definition of is $\varepsilon$-differential privacy, where a parameter $\varepsilon$, referred to as the privacy budget, is prespecified, such that a small $\varepsilon$ indicates a tight budget, associated with a stringent privacy protection protocol. Technically, $\varepsilon$-differential privacy requires that the alteration of any entry in the original dataset only leads to a small change of the output data’s distribution. This is quantified by the ratio of distributions before and after the alteration being bounded by $e^{\varepsilon}$.
On network data, however, the definition of one data entry may be construed differently (e.g., as a node or as an edge). There have been multiple ways to define differential privacy on network data, and a detailed review can be found in Jiang et al. (2021). Specifically, in this paper, we focus on node differential privacy (NDP), formally defined as below.
@@
$\mathbb{P}(\mathcal{M}(A)\in\Psi)\leq e^{\varepsilon}P(\mathcal{M}(A^{\prime})\in\Psi)$ (1)
@@ 3.3 General latent space models
We now proceed to introduce a class of statistical models for network data, referred to as the latent space model. We first introduce the design of our GRAND mechanism based on this model. And later we will demonstrate that our NDP guarantee still holds without the model.
The idea of latent space has been widely used in random networks, and was first formally introduced by Hoff et al. (2002). In this work, we adopt a slightly more general form than that in Hoff et al. (2002), though the model class is defined based on the same idea.
@@
We say that $A$ is a network generated from the general latent space model if there exists a distribution $F$ (unknown) on $\mathbb{R}^{d}$ and a known symmetric generative function $W:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1]$ such that $A$ can be generated through the procedure below:
$Z_{1},...,Z_{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F,\quad\quad A_{ij}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Bernoulli}(W(Z_{i},Z_{j})),\quad i>j.$
@@
We say that $A$ is a network generated from an inner-product latent space model if there exist distributions $F_{X}$ on $\mathbb{R}^{d}$ and $F_{\alpha}$ on $\mathbb{R}$, such that $A$ is generated as follows:
$(X_{1},\alpha_{1}),...,(X_{N},\alpha_{N})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F_{X}\times F_{\alpha},\quad\quad A_{ij}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Bernoulli}(W((X_{i},\alpha_{i}),(X_{j},\alpha_{j}))),\quad i>j$
where
$W((X_{i},\alpha_{i}),(X_{j},\alpha_{j}))=\sigma\big{(}X_{i}^{T}X_{j}+\alpha_{i}+\alpha_{j}\big{)}$ (2)
@@
If one is willing to introduce additional constraint bounds on the parameter space, the model can be further simplified as below.
Example 2 (Random dot product graph (RDPG) (Young and Scheinerman, 2007)).
We say $A$ is a network generated from a generalized random dot product model if there exists a distribution $F$ on $\mathbb{R}^{d}$, such that $A$ follows the generative procedure below [^2] There is an additional constraint on the domain of $F$ such that any two random vectors $Z_{1},Z_{2}$ from this distribution must have $Z_{1}^{T}Z_{2}\in[0,1]$.:
$Z_{1},...,Z_{N}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}F,\quad\quad A_{ij}\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\text{Bernoulli}(Z_{i}^{T}Z_{j}),\quad i>j$
where the support of $F$ guarantees the validity of the generalized inner product as a probability.
@@
Depending on the specific model within the latent space model family, several standard model estimation procedures are available to estimate $Z_{i}$’s with a given $A$. These include the gradient descent for the inner product latent space model (Ma et al., 2020) and the adjacency spectral embedding for the RDPG family (Athreya et al., 2018; Rubin-Delanchy et al., 2022).
@@
In most of the latent space models, the latent vectors are identifiable only up to some operator $\mathcal{O}^{d}$, depending on the specific format of $W$. For example, under the inner product latent space model or the RDPG, $\mathcal{O}^{d}$ can be any $d\times d$ orthogonal transformation. Such operators are usually not crucial when using these models, and would not result in difficulties in our analysis. For notational simplicity, throughout this paper, when we discuss the recovery of latent vectors or their distributions, the recovery can be up to such an unidentifiable operator.
Remark 3.
Another class of random network models, the graphon model (Bickel and Chen, 2009) can also be written in the form of the latent space model in Definition 2. However, there is one key difference: As stated in the definition, the latent space model typically assumes the link function $W$ to be known while the distribution of the latent vectors, $F$, can be unknown. In contrast, a graphon model has one dimensional latent variables $Z_{i}$’s with known $F$ (the uniform distribution), while the link function (graphon) $W$ is unknown. In this paper, we focus on the scenario of the latent space model with known $W$ but unknown $F$. We leave the study of node DP under the graphon model for future work.
Remark 4.
In the literature, some models introduce an additional parameter, which formulates the density of the network. However, the definition of this parameter differs by models; for instance, it appears as a scaling parameter in the RDPG and gRDPG frameworks (Athreya et al., 2021; Rubin-Delanchy et al., 2022) and as the intercept term in the inner product latent space model (Li et al., 2023a). To maintain a uniform format of our framework and focus only on the privacy aspect of networks, we will not introduce such a parameter. Our theory can be readily extended to incorporate such a parameter, if needed.

### v3

@@ Abstract
Differential privacy is a well-established framework for safeguarding sensitive information in data. While extensively applied across various domains, its application to network data — particularly at the node level — remains underexplored. Existing methods for node-level privacy either focus exclusively on query-based approaches, which restrict output to pre-specified network statistics, or fail to preserve key structural properties of the network. In this work, we propose GRAND (Graph Release with Assured Node Differential privacy), which is, to the best of our knowledge, the first network release mechanism that releases networks while ensuring node-level differential privacy and preserving structural properties. Under a broad class of latent space models, we show that the released network asymptotically follows the same distribution as the original network. The effectiveness of the approach is evaluated through extensive experiments on both synthetic and real-world datasets.
@@ 1 Introduction
Rapid technological advances have rendered complex networks pervasive — ranging from the Internet to international trade — yet these systems inherently present critical privacy challenges, exemplified by the need to protect identities in organ donation (Hizo-Abes et al., 2010; Marcus et al., 2023) and HIV transmission studies (Little et al., 2014; Abadie et al., 2021). Preemptively safeguarding such networks is uniquely difficult compared to conventional data sharing, as intricate structural dependencies, noise, and vast scales render simple mathematical models insufficient (Karwa et al., 2017; Hehir et al., 2022) while imposing rigorous computational and communication efficiency requirements on privacy-preserving algorithms (Guo et al., 2023).
@@
Differential privacy (DP; Dwork, 2006) has established itself as a standard for strict privacy protection, particularly for vulnerable outliers with extreme values. Mechanisms such as the Laplace (Dwork et al., 2006, 2014) and exponential (McSherry and Talwar, 2007) methods are widely employed in learning tasks (Lei et al., 2018; Soto et al., 2022; Lin et al., 2023; Xue et al., 2024; Ma et al., 2025). Furthermore, extensive research into DP variations—including relaxed (Abadi et al., 2016; Cai et al., 2019), local (Rohde and Steinberger, 2018), random (Hall et al., 2012), Kullback-Leibler (Wang et al., 2016), and Gaussian (Dong et al., 2022) frameworks—has facilitated broad application across academia and industry (Kaissis et al., 2020; Hie et al., 2018; Han et al., 2020; Santos-Lozada et al., 2020; Kenthapadi and Tran, 2018).
@@
Differential privacy in complex network data analysis has also been carefully explored (e.g., Karwa and Slavković, 2016; Rohde and Steinberger, 2018; Chang et al., 2024). However, as network data typically show structures distinct from that of Euclidean data, the preservation of differential privacy for network data entails unique challenges. First, differential privacy preservation may come at the cost of altering a network’s underlying utilities. This includes, but not limited to, crucial network properties, such as degree distributions, centrality and potentially community structures. Many existing methods may inadvertently alter the underlying distribution of edges and degrees. Meanwhile, edge-level modifications, such as edge addition or deletion, may also change node-level properties. In order to preserve a certain network property, therefore, many existing methods focus on statistic release rather than network release – that is, the release of certain summary statistics (e.g., degree distribution) with privacy guarantee, rather than an entire differentially private network that has all properties preserved.
@@
Second, the adoption of differential privacy may appear at different levels. On the one hand, many existing works consider edge differential privacy (Nissim et al., 2007), which protects privacy of individual edges. The adoption of edge differential privacy is natural and shows good theoretical properties (Mülle et al., 2015; Karwa et al., 2017; Fan et al., 2020; Hehir et al., 2022; Chang et al., 2024). On the other hand, the investigation and development of node differential privacy (Hay et al., 2009) is relatively rare. Node differential privacy guarantees that a released network should remain robust with the change of any node. However, the alteration of a node usually entails the alteration of all associated edges. The preservation of network properties under node differential privacy is therefore considered substantially more challenging (Kasiviswanathan et al., 2013; Hehir et al., 2022). In particular, the release of entire networks that follow node differential privacy has rarely been investigated in previous literature except for a few NP-hard theoretical algorithms (Borgs et al., 2015, 2018; Chen et al., 2024).
@@
In this work, we propose a novel mechanism, GRAND (Graph Release with Assured Node Differential privacy), to release networks that guarantees node differential privacy. Moreover, the proposed method is shown to preserve statistical network properties under the general latent space models, which include many widely used statistical network models as special cases. To the best of our knowledge, our method is the first practical approach that can (1) release a network (rather than a summary statistic), (2) ensure node differential privacy of the released network, and (3) is computationally feasible with utility preservation guarantees.
The paper proceeds as follows: Section 2 reviews existing literature and outlines our contributions, followed by formal definitions in Section 3. The GRAND mechanism and its theoretical properties are detailed in Sections 4 and 5, respectively. We evaluate performance via simulations in Section 6 and real-world applications in Section 7, concluding with discussions in Section 8.
2 Network Differential Privacy and Our Contributions
@@
Node differential privacy vs. edge differential privacy. Broadly speaking, existing works on network differential privacy can be categorized into two branches (Abawajy et al., 2016; Li et al., 2023c). One branch is edge differential privacy (EDP; Nissim et al., 2007). Intuitively, an EDP mechanism ensures that one cannot infer the pairwise relation between any two nodes in the network. Many methods have been proposed and discussed regarding the protection of EDP, which protect the privacy of individual edges. This can be achieved through mechanisms such as randomized responses (Mülle et al., 2015; Hehir et al., 2022), method of moments (Chang et al., 2024), or noise injection (Karwa and Slavković, 2016; Fan et al., 2020). The other branch is node differential privacy (NDP; Hay et al., 2009). An NDP mechanism (to be formally defined later) ensures that one cannot infer an individual node in the network.
@@
Notice that, a node’s information includes its pairwise relations to other nodes, and hence node differential privacy is more strict than edge differential privacy. As a result, node-level privacy protection is considerably more challenging than edge-level privacy. Applying edge-level DP methods to node-level settings usually leads to degenerate results (Kasiviswanathan et al., 2013; Hehir et al., 2022; Chang et al., 2024).
@@
However, we advocate that preserving node differential privacy can be critical in many scenarios, because nodes can represent individual users, customers, and patients, whose identities are sensitive and vulnerable to de-anonymization. Encoding node differential privacy, while preserving the underlying network’s properties is the key motivation of this paper.
@@
Network release vs. statistic release. The adoption of differential privacy in network analysis, either at the node level or at the edge level, involves introducing random noise, which may lead to unintended consequences. First, existing differentially private mechanisms that are designed for Euclidean data may inadvertently change the underlying structure or properties of networks. For example, adding or deleting edges may lead to the alteration of the degree distribution, transitivity, and the number of components. In general, releasing a differentially private network while preserving all original properties can be very challenging.
@@
With the aforementioned limitations, it is then evident that directly releasing a network would be substantially more flexible for subsequent use. Ideally, if the released network is almost the same as the original network, known as an informative release (Wasserman and Zhou, 2010), we can use the privatized network in an arbitrary way in downstream tasks. In particular, when the released network has most network properties maintained the same, it essentially subsumes most statistic release. Along this path, many have explored the novel ways of generating synthetic networks under edge differential privacy such as Karwa et al. (2017); Qin et al. (2017); Hehir et al. (2022); Guo et al. (2023); Chang et al. (2024). For node differential privacy, the only known methods in the literature (Borgs et al., 2015, 2018; Chen et al., 2024) are NP-hard, thus are only of theoretical interest.
Moving beyond the prioritization of network release strategies, we also consider the use of data-perturbation mechanisms (as opposed to synthetic network generation). Data-perturbation mechanisms inject noise into the original network. A key feature is that it retains a bijective (one-to-one) correspondence between the anonymized nodes and the original individuals. This avoids the construction of completely artificial nodes as used in synthetic networks, giving substantial advantages with respect to data expandability, decentralizability and robustness (details in Section 8). For this reason, we explicitly prioritize the data-perturbation paradigm.
Our contributions. The preceding discussion provides the motivation for a data-perturbation mechanism that releases networks (rather than statistics) with node-level differential privacy and guarantees of preserving network structures. This is expected to be a crucial tool in practice, while such a tool is not yet available in the literature. In this paper, we introduce a computationally feasible method that achieves all the above objectives simultaneously. To the best of our knowledge, we are among the first to achieve this goal in the literature.
@@ 3 Node Differential Privacy and Latent Space Models
Notations. For any positive integer $K$, we denote the set $\{1,\ldots,K\}$ by $[K]$. Consider a network of $N$ individuals (nodes), where the edges represent interactions or relationships between nodes, such as friendship, collaboration, or transactions (Newman, 2018). Denote the node set by $[N]$. We assume the network is unweighted and undirected. The network can be uniquely represented by an $N\times N$ adjacency matrix $A$ where $A_{ij}=1$ if individuals $i$ and $j$ are connected, and $A_{ij}=0$ otherwise. Denote the node and edge sets of network $A$ by $\mathcal{V}(A)$ and $\mathcal{E}(A)$, respectively. For any vector $z\in\mathbb{R}^{d}$, we use $\|{z}\|$ to denote its Euclidean norm.
@@ 3.1 Node Differential Privacy
Suppose that the network information is sensitive (e.g., a cancer patient network) and many nodes in $A$ may not be willing to release their identity or even their presence in the network. Private information may still be vulnerable to leakage if one simply releases the network to the public, even with de-anonymization (Narayanan and Shmatikov, 2008). In many cases, one may infer a node’s identity according to certain connectivity patterns of a de-anonymized network. Alternatively, a node’s identity may be revealed due to adversarial attacks.
We adopt differential privacy (Dwork, 2006) as the standard for privacy protection. One commonly used definition of is $\varepsilon$-differential privacy, where a parameter $\varepsilon$, referred to as the privacy budget, is prespecified, such that a small $\varepsilon$ indicates a tight budget, associated with a stringent privacy protection protocol. Technically, $\varepsilon$-differential privacy requires that the alteration of any entry in the original dataset only leads to a small change of the output data’s distribution, quantified by the ratio of distributions before and after the alteration being bounded by $e^{\varepsilon}$.
On network data, however, the definition of one data entry may be construed differently (e.g., as a node or as an edge). There have been multiple ways to define differential privacy on network data, and a detailed review can be found in Jiang et al. (2021). Specifically, we focus on node differential privacy (NDP), formally defined as below.
@@
$\mathbb{P}(\mathcal{M}(A)\in\Psi)\leq e^{\varepsilon}\mathbb{P}(\mathcal{M}(A^{\prime})\in\Psi)$ (1)
@@ 3.2 General Latent Space Models
We now proceed to introduce a class of statistical models for network data, referred to as the latent space model. We first introduce the design of our GRAND mechanism based on this model. And later we will demonstrate that our NDP guarantee still holds without the model. The idea of latent space has been widely used in random networks, and was first formally introduced by Hoff et al. (2002).
@@
We say that $A$ is a network generated from the general latent space model if there exists a distribution $F$ (unknown) on $\mathbb{R}^{d}$ and a known symmetric generative function $W:\mathbb{R}^{d}\times\mathbb{R}^{d}\to[0,1]$ such that $A$ can be generated as below:
$Z_{1},\ldots,Z_{N}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F,\quad\quad A_{ij}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\text{Bernoulli}(W(Z_{i},Z_{j})),\quad i>j.$
@@
We say that $A$ is a network generated from an inner product latent space model if there exist distributions $F_{X}$ on $\mathbb{R}^{d}$ and $F_{\alpha}$ on $\mathbb{R}$, such that $A$ is generated as follows:
$(X_{1},\alpha_{1}),\ldots,(X_{N},\alpha_{N})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F_{X}\times F_{\alpha},\quad\quad A_{ij}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\text{Bernoulli}(W((X_{i},\alpha_{i}),(X_{j},\alpha_{j}))),\quad i>j,$
$W((X_{i},\alpha_{i}),(X_{j},\alpha_{j}))=\sigma(X_{i}^{\top}X_{j}+\alpha_{i}+\alpha_{j})$ (2)
@@
Example 2 (Random Dot Product Graph (Young and Scheinerman, 2007)).
We say $A$ is a network generated from a generalized random dot product graph (RDPG) if there exists a distribution $F$ on $\mathbb{R}^{d}$, such that $A$ follows the generative procedure below:
$Z_{1},\ldots,Z_{N}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}F,\quad\quad A_{ij}\stackrel{{\scriptstyle\mathrm{ind.}}}{{\sim}}\text{Bernoulli}(Z_{i}^{\top}Z_{j}),\quad i>j,$
where the support of $F$ guarantees that the generalized inner product is a probability.
@@
In most of the latent space models, the latent vectors are identifiable only up to some operator $\mathcal{O}^{d}$, depending on the specific format of $W$. For example, under the inner product latent space model or the RDPG, $\mathcal{O}^{d}$ can be any $d\times d$ orthogonal transformation. Such operators are usually not crucial when using these models, and would not result in difficulties in our analysis. For notational simplicity, in this paper, when we discuss the recovery of latent vectors or their distributions, the recovery can be up to such an unidentifiable operator.
@@
In the literature, some models introduce an additional parameter, which formulates the density of the network. However, the definition of this parameter differs by models; for instance, it appears as a scaling parameter in the RDPG and gRDPG frameworks (Athreya et al., 2021; Rubin-Delanchy et al., 2022) and as the intercept term in the inner product latent space model (Li et al., 2023a). To maintain a uniform format of our framework and focus only on the privacy aspect of networks, we will not introduce such a parameter. Our theory can be readily extended to incorporate such a parameter and we will briefly mention the corresponding minor changes in Section 5.